OpenStudy (anonymous):

Find all polar coordinates of point P where P = (5,-pi/6)

OpenStudy (dumbcow):

x = r*cos y = r*sin y/x = tan x^2 + y^2 = r^2

OpenStudy (dumbcow):

ha i totally misunderstood, i didn't see they already gave you a point in polar form  this gives 4 possible coordinates which all end up at same point P (5, -pi/6) (5, 11pi/6) (-5, -7pi/6) (-5, 5pi/6)